# Adaptive Multilevel Splitting: First Application to Rare-Event Derivative Pricing

### v1

@@ 4.3.3 Analysis of option pricing results
Role of the importance function. Performance depends only moderately on $\xi$. Both tested choices are effective; the Black–Scholes-based function (AMS2) provides smoother guidance via the $\Phi(d_{2})$ term, improving sampling efficiency in some cases.
Consistency across option types. Efficiency gains hold for European and Asian binaries under both models, indicating robustness across payoffs and dynamics.
Overall, AMS delivers unbiased estimates with significant computational savings relative to Monte Carlo, and remains competitive with advanced variance reduction methods, particularly in rare-event regimes.
4.4 Extreme case analysis
@@
AMS attains the target within 30 seconds, confirming its robustness in extreme rare-event regimes where standard Monte Carlo is computationally infeasible.
@@ 5.1 Conclusions
This study establishes adaptive multilevel splitting (AMS) as a computationally superior method for pricing binary options in rare-event regimes. Across both Black–Scholes and Heston models, AMS achieves speedups of up to 200 over standard Monte Carlo while maintaining unbiasedness, and consistently outperforms variance-reduction baselines such as MLMC and antithetic variates.
To our knowledge, this is the first application of AMS to financial rare-event pricing. Benchmarking against the closest variance reduction methods in finance confirms its superior efficiency in deep out-of-the-money regimes, where conventional techniques become computationally infeasible.
The practical implications are significant: AMS renders previously intractable problems feasible, enabling tighter spreads and deeper liquidity for rare-event derivatives, with direct relevance for parametric insurance and catastrophe-linked products.
@@ 5.2 Future developments
A first direction is risk management, where AMS could improve the computation of tail risk measures. Value-at-Risk (VaR), defined as the loss threshold exceeded with small probability, is a rare-event problem. Existing Monte Carlo and importance sampling approaches are widely used [45, 46]; AMS offers the potential for more accurate and efficient estimates, relevant for stress testing and regulatory capital.
A second extension concerns model coverage. Incorporating exotic payoffs and multi-asset structures would broaden applicability, enabling AMS to address higher-dimensional rare-event problems and increasing the versatility of the package for quantitative finance.
A third avenue is methodological. Rough volatility models such as Bergomi [47] pose challenges because fractional Brownian motion violates the Markov property central to AMS. One possible solution is a lifted Markovian approximation embedding the non-Markovian dynamics in higher-dimensional state space [48], potentially extending AMS to this class of models.
@@ Appendix A C++ implementation with R interface via Rcpp
No R package currently provides AMS functionality for financial applications. To fill this gap, a dedicated implementation was developed in C++ [49] with an R interface via Rcpp [50].
@@
The Rcpp interface exposes all algorithmic parameters and diagnostics within the R environment, while computationally intensive tasks remain in C++. This design combines the usability of R with near-native performance, enabling practical deployment of AMS in quantitative finance.
@@ A.1 Core implementation
payoff Evaluates six exotic payoffs: digital call, digital put, asian digital call, asian digital put, lookback call, and lookback put.
@@
Lookback options are implemented but excluded from the numerical study, as discretization error under discrete monitoring [51] prevents reliable benchmarking. Experiments are restricted to European and Asian binaries.
@@ Appendix B Tables underlying the figures
Table 6: Computational times (in seconds) for different relative accuracy levels in the Heston digital put experiment.
@@
0.20 85.51 75.2 54.82 0.8 0.56
0.15 144.83 122.11 115.2 1.21 1.12
0.10 307.55 231.21 188.73 1.67 1.71
0.05 1235.77 910.3 659.9 5.55 6.09
Table 7: Computational times (in seconds) for different relative accuracy levels in the Black–Scholes Asian digital call experiment.
Relative accuracy MC MCA MLMC AMS 1 AMS 2
0.20 27.53 26.23 15.38 1.16 0.2
0.15 49.85 46.84 29.62 2.35 0.44
0.10 105.52 92.62 39.69 3.85 0.79
0.05 415.65 305.63 228.73 9.8 15.27
Table 8: Computational times (in seconds) for different relative accuracy levels in the Black–Scholes Asian digital put experiment.
Relative accuracy MC MCA MLMC AMS 1 AMS 2
0.20 28.11 26.52 16.7 1.351 0.316
0.15 50.73 47.41 27.91 2.160 0.504
0.10 106.75 95.31 38.97 3.910 0.815
0.05 421.29 309.77 214.84 9.99 14.11
@@
Relative accuracy MC MCA MLMC AMS 1 AMS 2
0.20 82.22 73.38 53.6 4.3 2.62
0.15 139.31 117.6 108.37 10.94 6.76
0.10 301.53 228.35 185.88 14.08 10.794
0.05 1221.18 899.04 640.85 106.05 36.56
Table 10: Computational times (in seconds) for different relative accuracy levels in the Heston Asian digital put experiment.
Relative accuracy MC MCA MLMC AMS 1 AMS 2
0.20 84.06 74.19 53.8 4.8 2.74
0.15 141.35 121.7 109.32 11.24 6.55
0.10 312.04 230.14 186.05 15.28 11
0.05 1212.5 901.29 645.01 110.6 35.41

### v2

@@ 4.4 Analysis of option pricing results
Role of the importance function. The two importance functions produce curves that are close in Figures 3 and 4, indicating that AMS is not overly sensitive to this modelling choice in the configurations tested. Nevertheless, the Black-Scholes based importance function (AMS2) tends to yield slightly lower times at a fixed relative accuracy in several panels, consistent with the intuition that the $\Phi(d_{2})$ term provides a smoother and more informative guidance towards the rare-event region.
Consistency across option types. The patterns observed in Figure 3 for the digital and up-and-in barrier options are essentially reproduced in Figure 4 for the Asian digital case: the relative ordering of the methods remains unchanged, and AMS consistently emerges as the most efficient approach in all deep out-of-the-money configurations, for both European and Asian binaries and under both Black–Scholes and Heston dynamics. This indicates that the efficiency gains achieved by AMS are robust with respect to the payoff specification as well as to the underlying volatility model.
Taken together, Figures 3 and 4 demonstrate that AMS produces unbiased estimates while delivering substantial computational savings relative to standard Monte Carlo, and remains competitive against advanced variance–reduction techniques such as MLMC or control variates. The advantage is particularly pronounced in the rare-event regime that motivates the use of AMS, where conventional methods become prohibitively expensive.
4.4.1 Extreme case analysis
@@
AMS attains the target within 30 seconds, confirming its robustness in extreme rare event regimes where standard Monte Carlo is computationally infeasible.
4.4.2 Multi-asset digital option analysis
To illustrate the flexibility of AMS in a multidimensional setting, we consider a synthetic three-asset digital option that does not correspond to a traded product and is not included in the amsSim library. The contract is introduced solely as a stress test in a multi-asset rare-event scenario.
We work under a Black–Scholes framework with three underlying assets $S^{(1)},S^{(2)},S^{(3)}$ following correlated geometric Brownian motions under the risk–neutral measure $\mathbb{Q}$. The driving Brownian motions have constant pairwise correlation $\rho=0.2$. In the numerical experiment we take a common initial price $S_{0}^{(i)}=1$ for all three assets, a constant volatility $\sigma=0.2$, and a constant risk-free rate $r=0.03$.
At maturity $T$ we define the arithmetic average
$\bar{S}_{T}=\frac{1}{3}\bigl(S_{T}^{(1)}+S_{T}^{(2)}+S_{T}^{(3)}\bigr)$
and the maximum pairwise terminal price difference
$\Delta_{\max}=\max\Bigl\{\bigl|S_{T}^{(1)}-S_{T}^{(2)}\bigr|,\bigl|S_{T}^{(2)}-S_{T}^{(3)}\bigr|,\bigl|S_{T}^{(1)}-S_{T}^{(3)}\bigr|\Bigr\}.$
Given a dispersion threshold $L>0$ and an average level $K_{\text{avg}}>0$, the payoff is
$\Pi=\begin{cases}1,&\text{if }\Delta_{\max}>L\text{ and }\bar{S}_{T}>K_{\text{avg}},\\[4.0pt] 0,&\text{otherwise,}\end{cases}$ (7)
In the parameter setting used here, the maximum price difference must exceed $L=1$, while the arithmetic average must satisfy $\bar{S}_{T}>K_{\text{avg}}=1.4$, with $S_{0}^{(i)}=1$ for all assets.
The payoff is path-independent but strongly non-linear and multidimensional, and no closed-form solution is available in the classical Black–Scholes setting.
The payoff is deliberately designed to combine two competing features: on the one hand, it rewards configurations in which the assets exhibit strong cross-sectional dispersion (two prices moving far above the strike while the remaining one moves far below it), so that the rare event is driven by extreme relative movements; on the other hand, an additional condition on the cross-sectional average imposes a common level constraint that prevents trivial scenarios where the whole basket drifts to a very low mean.
For the present three-asset option we define the score function as
$\xi=\Delta_{\max}+\bar{S}_{T},$
so that trajectories closer to the rare-event region $\{\Delta_{\max}>L,\;\bar{S}_{T}>K_{\text{avg}}\}$ receive larger scores. Choosing $L_{\max}=L+K_{\text{avg}}$ such that
$(\Delta_{\max},\bar{S}_{T})\in\{\Delta_{\max}>L,\;\bar{S}_{T}>K_{\text{avg}}\}\;\Rightarrow\;\xi\geq L_{\max}$
ensures that the sufficient condition of [brehier2016unbiasedness, Theorem 3.2] holds and the AMS estimator remains unbiased.
In the multi-asset setting, the set of viable benchmarks becomes considerably more restricted. MLMC, although effective for one-dimensional path functionals, does not admit a straightforward extension to this payoff, whose rare-event structure combines cross-sectional dispersion and joint level constraints; constructing a suitable hierarchy and coupling mechanism is non-trivial.
Importance sampling (IS) is even more delicate. In high dimensions, an effective change of measure must simultaneously bias the process towards large relative deviations and a higher average level, leading to a highly anisotropic rare set. Identifying such a tilt is problem-specific and unstable. Even cross-entropy methods, often used to automate the search for an optimal IS density, become difficult to apply: selecting a parametric family capable of capturing the joint tail behaviour is non-obvious.
Due to these challenges, we restrict the comparison to standard Monte Carlo and Monte Carlo with antithetic variates.
The corresponding performance metrics, together with those of the AMS estimator, are reported in Table 4.
Table 4: Computational times (in seconds) for different relative accuracy levels in the multi-dimension call experiment.
Relative accuracy MC MCA AMS
0.05 23.62 19.22 0.44
0.03 67.89 58.93 1.25
0.01 555.49 438.13 12.19
As in the one-dimensional setting, AMS achieves a substantial computational cost reduction for comparable variance, confirming its effectiveness also in this multidimensional rare-event scenario.
@@ 5.1 Conclusions
This study establishes adaptive multilevel splitting (AMS) as a computationally superior method for pricing binary options in rare event regimes. Across both Black–Scholes and Heston models, AMS achieves speedups of up to 200 over standard Monte Carlo while maintaining unbiasedness, and consistently outperforms variance reduction baselines such as MLMC and antithetic variates.
To the best of our knowledge, this is the first application of AMS to financial rare event pricing. Benchmarking against the closest variance reduction methods in finance confirms its superior efficiency in deep out-of-the-money regimes, where conventional techniques become computationally infeasible.
The practical implications are significant: AMS renders previously intractable problems feasible, enabling tighter spreads and deeper liquidity for rare event derivatives, with direct relevance for parametric insurance and catastrophe-linked products.
@@ 5.2 Future developments
A first direction is risk management, where AMS could improve the computation of tail risk measures. Value-at-Risk (VaR), defined as the loss threshold exceeded with small probability, is a rare event problem. Existing Monte Carlo and importance sampling approaches are widely used [hong2014monte, sun2010asymptotic]; AMS offers the potential for more accurate and efficient estimates, relevant for stress testing and regulatory capital.
A second extension concerns model coverage. Incorporating exotic payoffs and multi-asset structures would broaden applicability, enabling AMS to address higher-dimensional rare event problems and increasing the versatility of the package for quantitative finance.
A third avenue is methodological. Rough volatility models such as Bergomi [gatheral2017rough] pose challenges because fractional Brownian motion violates the Markov property central to AMS. One possible solution is a lifted Markovian approximation embedding the non-Markovian dynamics in a higher-dimensional state space [zhu2021markovian], potentially extending AMS to this class of models.
@@ Appendix A C++ implementation with R interface via Rcpp
Prior to this work, no R package provided AMS functionality for financial applications. To fill this gap, a dedicated implementation was developed in C++ [press2007numerical] with an R interface via Rcpp [eddelbuettel2013seamless].
@@
The Rcpp interface exposes all algorithmic parameters and diagnostics within the R environment, while computationally intensive tasks remain in C++. This design combines the usability of R with near native performance, enabling practical deployment of AMS in quantitative finance. The complete implementation is distributed as the R package amsSim [mio], available on CRAN.
@@ A.1 Core implementation
payoff Evaluates six exotic payoffs: digital call, digital put, Asian digital call, Asian digital put, up-and-in barrier call, and up-and-in barrier put.
@@ Appendix B Tables underlying the figures
Table 7: Computational times (in seconds) for different relative accuracy levels in the Heston digital up-and-in barrier call experiment.
@@
0.20 83.24 75.94 54.12 0.64 0.65
0.15 148.98 119.32 116.35 1.34 1.62
0.10 322.72 238.31 196.04 4.25 3.23
0.05 1311.89 962.51 699.78 11.36 13.38
Table 8: Computational times (in seconds) for different relative accuracy levels in the Black-Scholes Asian digital call experiment.
Relative accuracy MC MCA MCV MLMC AMS 1 AMS 2
0.20 27.53 26.23 21.77 15.38 1.16 0.2
0.15 49.85 46.84 40.11 29.62 2.35 0.44
0.10 105.52 92.62 94.62 39.69 3.85 0.79
0.05 415.65 305.63 309.81 228.73 9.8 15.27
@@
Relative accuracy MC MCA MCV MLMC AMS 1 AMS 2
0.20 82.22 73.38 34.31 53.6 4.3 2.62
0.15 139.31 117.6 60.86 108.37 10.94 6.76
0.10 301.53 228.35 116.19 185.88 14.08 10.79
0.05 1221.18 899.04 474.85 640.85 106.05 36.56